The Adaptive DD Preprocessor ADDPre and Its Use in the Parallel Code FEM^{◯◯}BEM

M. Kuhn , G. Haase, B. Heise, U. Langer Institute of Mathematics Johannes Kepler University Linz Austria

ABSTRACT. Applying Domain Decomposition (DD) methods for solving partial differential equations approximately via a finite element (f.e.), boundary element (b.e.) or a mixed f.e./b.e. discretization requires the partitioning of the data. The Adaptive Domain Decomposition Preprocessor ADDPre realizes automatically such a decomposition of the domain under consideration into p subdomains, where p is the number of processors to be used. Three codes Decomp, Tri2DD and AdapMesh which are based on the recursive spectral bisection method, a transformation of triangular-based data into a special DD-data format and a parallel mesh generator, respectively, are involved in the preprocessing. We discuss the (iterative) procedure of the data partitioning aiming at a well load-balanced decomposition which allows also the grids to be adapted to singularities. The resulting decomposition is being described using the DD-data format, which is then the input for the parallel code FEM^{◯ ◯}BEM .

Several examples, including electric motors, and numerical results obtained with up to 64 processors will demonstrate the performance of both, the preprocessor ADDPre and the code FEM^{◯ ◯}BEM .